OpenStudy (anonymous):

A 10-gallon container of water costs (y+2) dollars while a 5-gallon container of water costs (y-2) dollars. If the price per gallon of water is the same for both containers, the equation below can be used to solve for y. y+2/10 = y_2/5. What is the value, in dollars, of y?

OpenStudy (hoblos):

y+2/10 = y-2/5 cross multiply.. 5y+10 = 10y-20 30 = 5y y=6

hero (hero):